# A Galactic Transformation—Understanding the SMC’s Structural and Kinematic Disequilibrium

### v1

@@ 4.5 Limitations and Future Work
A limitation with the B12 simulation is the limited DM mass resolution ($10^{6}$ M per particle). However, as shown in Rathore2025b, this is not a concern for understanding the morphological and kinematic transformation of the Clouds driven by their recent collision.
@@
The LMC’s galactocentric velocity at “present day” in the B12 simulation has been matched to the Hubble Space Telescope (HST) PMs from (Kallivayalil2006; Kallivayalil2006b). However, the observed LMC PMs have since been updated with third epoch HST observations (Kallivayalil2013), and further refined by Gaia (Luri2021). The galactocentric position and velocity of the SMC at present day are also not quite correct in the B12 simulation. Further, the maximum post-collision SMC-LMC separation in the simulation is $\approx 10$ kpc, whereas the observed 3D separation between the Clouds is $\approx 20$ kpc. The inability of the simulation to fully match the observed position and velocity of the Clouds limits direct comparisons with observations. However, direct comparison with observations is not the goal of our study. Rather, as mentioned in section 2.1, our goal is to understand the physical principles that operate to transform the SMC’s structure and kinematics due to the recent SMC-LMC collision, and guide the interpretation of the observational data. We also refrain from using the word “present day” for labeling different epochs in the simulation, and rather perform a generic study of the SMC’s internal state as a function of time elapsed since the collision.
@@
The B12 simulation includes neither supernova feedback nor a MW/LMC CGM. Both of these factors can significantly affect the gas morphology and kinematics of the SMC (Lucchini2021). Given the SMC’s high star-formation rate (Harris2006; Rubele2018; Massana2022), feedback is expected to be significant over the past $\sim 100$ Myr. However, neither feedback nor ram pressure from the LMC/MW CGM will enable the SMC to retain its gas or support the persistence of internal rotation. As such, the inclusion of these physical effects would only strengthen the arguments made in this paper. Moreover, we have demonstrated that the ram pressure exerted by the LMC’s gas disk on the SMC is significantly stronger than that expected by the LMC or MW CGM.
@@
There is a large amount of observational evidence (mostly) related to the LMC’s disturbed morphology and kinematics that suggests a recent ($\sim 100$ Myr ago) close encounter ($\sim 2$ kpc) between the Clouds (see section 1). However, the SMC-LMC orbit is not tightly constrained and a collision, while allowed within the proper motion error space, is not the mean result. Different choices of the SMC center (see right panel in Figure 1) result in differences in the SMC 3D velocity measurements reported using HST (Zivick2018) and Gaia (Luri2021). In this work, we argue that the stellar photometric center, rather than the HI kinematic center should be used to determine the 3D velocity of the SMC. However, it remains to be tested how a collision impacts the measurements of the SMC’s stellar proper motions compared to the center of mass motion of the SMC’s DM.
@@
Further, it is likely that the DM halos of the LMC and SMC are significantly deformed due to their interactions with each other (H. Foote et al. 2025 in prep) and interactions with the MW (GC2019). Thus, it is likely that analytical orbits are not able to capture the full complexity of the SMC-LMC interaction and need to be significantly revised using time-evolving potentials.
@@ 5 Conclusions
    The SMC’s tidal tail that formed from a direct SMC-LMC collision can explain the SMC’s large LoS depth and stellar axis ratio: After the SMC-LMC collision, the LMC’s tides cause a break in the SMC’s stellar density profile at $R<4$ kpc (Figure 5) and a tidal tail and bridge are formed. The stellar tidal tail explains the observed SMC’s large axis ratio (>4:1>4:1, Figure 6), which is not achieved without a direct collision. The stellar LoS depth of the SMC is naturally explained if the tidal tail is oriented along the LoS (Figure 7 right panel) (see also, DB2012) . Further, the tidal elongation has a significant component perpendicular to the SMC’s plane of rotation (Figure 8), and the orientation of the tail cannot be used to constrain the SMC’s inclination. The formation of an SMC tidal tail is a generic prediction for all tidal SMC-LMC simulations (e.g. Connors2006). We have demonstrated that a direct collision can also create this morphology. Further, the gas tidal tail born out of the SMC-LMC collision can in-part explain the existence of gas components at different LoS distances in the observed SMC (Figure 14).
@@
    The SMC-LMC collision converts the SMC from a rotationally supported disk galaxy to a dispersion dominated galaxy: The pre-collision simulated SMC has a coherent stellar rotation field by design. Post-collision, the simulated SMC’s stars at $R>2$ kpc exhibit radial outward motions (Figure 9), with a small (<10<10 km s-1) remnant rotation in the inner $R<2$ kpc. The post-collision simulated SMC’s internal stellar velocity field is consistent with observations (Niederhofer2021; Zivick2021). Further, the SMC’s stellar velocity field in the scenario where the Clouds do not collide, remains a rotating disk (peak velocity $\sim 40$ km s-1). This is inconsistent with observations, indicating that strong LMC tides (like in a direct collision) are needed to sufficiently disrupt the SMC’s internal velocity field. The SMC’s rotation to dispersion ratio significantly reduces post-collision (<0.4<0.4, Figure 10), indicating the SMC is in the midst of a transformation from a rotationally supported disk to a dispersion dominated spheroid.
@@
    The SMC’s gas kinematics is dominated by radially outward motions as opposed to rotation: Pre-collision, the simulated SMC has a coherent gas rotation field, by design. Post-collision, the simulated SMC’s gas also exhibits radially outward motions and the rotation field is destroyed (Figure 11). The resulting diverging velocity field manifests as a velocity gradient (408040-80 km s-1) when viewed at an inclination. Hence, the SMC’s observed HI velocity gradient is likely due to radially outward motions of gas viewed at an inclination, and is not a signature of rotation. This finding is consistent with the kinematics of the SMC’s young stellar populations (<100<100 Myr, Murray2019; Nakano2025a; Nakano2025b). The SMC’s gas kinematics cannot be used to place the SMC on the BTFR (Figure 1 left panel).
@@
Our work highlights the crucial role of galactic collisions in transforming dIrr type galaxies to dE/dSph type over short timescales ($\sim 100$ Myr). Further, ram pressure exerted by the primary galaxy’s ISM during such close encounters can provide a significant velocity kick to the secondary’s gas, assisting the removal of gas and contributing to quenching of low mass galaxies. The SMC gives a front row view of group processes driving dramatic morphological and kinematic transformations, without requiring the presence of sustained environmental effects (like a cluster ICM). Such group processing may explain the how dEs can exist around massive galaxies like M31.
@@
Himansh Rathore would like to thank Annapurni Subramaniam, Smitha Subramanian, Claire Murray, Paul Zivick, Andrew Fox, Ivanna Escala, Nico Garavito Camargo, Kathryne J. Daniel, S.R. Dhanush, Mathieu Renzo, Yumi Choi, Dennis Zaritsky, Mariarosa Cioni, Lara Cullinane and Hayden Foote for interesting scientific discussions. Himansh Rathore and Gurtina Besla are supported by NASA FINESST 80NSSC24K1469, NASA ATP 80NSSC24K1225 and NSF CAREER AST 1941096. This work utilized the Puma and ElGato High Performance Computing clusters at the University of Arizona. The Theoretical Astrophysics Program (TAP) at the University of Arizona provided resources to support this work. We respectfully acknowledge the University of Arizona is on the land and territories of Indigenous peoples. Today, Arizona is home to 22 federally recognized tribes, with Tucson being home to the O’odham and the Yaqui. The university strives to build sustainable relationships with sovereign Native Nations and Indigenous communities through education offerings, partnerships, and community service.

### v2

@@ IV.5 Limitations and Future Work
A limitation with the B12 simulation is the limited DM mass resolution ($10^{6}$ M per particle). However, as shown in Rathore et al. (2025a), this is not a concern for understanding the morphological and kinematic transformation of the Clouds driven by their recent collision.
@@
The LMC’s galactocentric velocity at “present day” in the B12 simulation has been matched to the Hubble Space Telescope (HST) PMs from (Kallivayalil et al., 2006b, a). However, the observed LMC PMs have since been updated with third epoch HST observations (Kallivayalil et al., 2013), and further refined by Gaia (Gaia Collaboration et al., 2021). The galactocentric position and velocity of the SMC at present day are also not quite correct in the B12 simulation. Further, the maximum post-collision SMC-LMC separation in the simulation is $\approx 10$ kpc, whereas the observed 3D separation between the Clouds is $\approx 20$ kpc. The inability of the simulation to fully match the observed position and velocity of the Clouds limits direct comparisons with observations. However, direct comparison with observations is not the goal of our study. Rather, as mentioned in section II.1, our goal is to understand the physical principles that operate to transform the SMC’s structure and kinematics due to the recent SMC-LMC collision, and guide the interpretation of the observational data. We also refrain from using the word “present day” for labeling different epochs in the simulation, and rather perform a generic study of the SMC’s internal state as a function of time elapsed since the collision.
@@
The B12 simulation includes neither supernova feedback nor a MW/LMC CGM nor reionization heating of the Clouds’ gas. All of these factors can significantly affect the gas morphology and kinematics of the SMC (Lucchini et al., 2021). Given the SMC’s high star-formation rate (Harris and Zaritsky, 2006; Rubele et al., 2018; Massana et al., 2022), feedback is expected to be significant over the past $\sim 100$ Myr. However, neither feedback nor ram pressure from the LMC/MW CGM nor reionization heating will enable the SMC to retain its gas or support the persistence of internal rotation (see also Gatto et al. 2013; Mayer et al. 2006). As such, the inclusion of these physical effects would only strengthen the arguments made in this paper. Moreover, we have demonstrated that the ram pressure exerted by the LMC’s gas disk on the SMC is significantly stronger than that expected by the LMC or MW CGM.
@@
There is a large amount of observational evidence (mostly) related to the LMC’s disturbed morphology and kinematics that suggests a recent ($\sim 100$ Myr ago) close encounter ($\sim 2$ kpc) between the Clouds (see section I). However, the SMC-LMC orbit is not tightly constrained and a collision, while allowed within the proper motion error space, is not the mean result. Different choices of the SMC center (see right panel in Figure 1) result in differences in the SMC 3D velocity measurements reported using HST (Zivick et al., 2018) and Gaia (Gaia Collaboration et al., 2021). In this work, we argue that the stellar photometric center, rather than the HI kinematic center should be used to determine the 3D velocity of the SMC. However, it remains to be tested how a collision impacts the measurements of the SMC’s stellar proper motions compared to the center of mass motion of the SMC’s DM.
@@
Further, it is likely that the DM halos of the LMC and SMC are significantly deformed due to their interactions with each other (H. Foote et al. 2025 in prep) and interactions with the MW (Garavito-Camargo et al., 2019). Thus, it is likely that analytical orbits are not able to capture the full complexity of the SMC-LMC interaction and need to be significantly revised using time-evolving potentials (Foote et al., 2026).
Finally, the present work primarily focusses on the SMC’s main body ($R<4$ kpc), and not the outskirts ($R>4$ kpc). The SMC’s outskirts harbor several interesting tidal features like the “SMC wing” (see Almeida et al. 2024 and references therein, Escala et al. 2026 in prep), which could also provide important constraints on the LMC-SMC interactions. This will be a subject of future analyses.
@@ V Conclusions
    The SMC’s tidal tail that formed from a direct SMC-LMC collision can explain the SMC’s large LoS depth and stellar axis ratio: After the SMC-LMC collision, the LMC’s tides cause a break in the SMC’s stellar density profile at $R<4$ kpc (Figure 5) and a tidal tail and bridge are formed. The stellar tidal tail explains the observed SMC’s large axis ratio (>4:1>4:1, Figure 6), which is not achieved without a direct collision. The stellar LoS depth of the SMC is naturally explained if the tidal tail is oriented along the LoS (Figure 7 right panel) (see also, Diaz and Bekki, 2012) . Further, the tidal elongation has a significant component perpendicular to the SMC’s plane of rotation (Figure 8), and the orientation of the tail cannot be used to constrain the SMC’s inclination. The formation of an SMC tidal tail is a generic prediction for all tidal SMC-LMC simulations (e.g. Connors et al., 2006). We have demonstrated that a direct collision can also create this morphology. Further, the gas tidal tail born out of the SMC-LMC collision can in-part explain the existence of gas components at different LoS distances in the observed SMC (Figure 14).
@@
    The SMC-LMC collision converts the SMC from a rotationally supported disk galaxy to a dispersion dominated galaxy: The pre-collision simulated SMC has a coherent stellar rotation field by design. Post-collision, the simulated SMC’s stars at $R>2$ kpc exhibit radial outward motions (Figure 9), with a small (<10<10 km s-1) remnant rotation in the inner $R<2$ kpc. The post-collision simulated SMC’s internal stellar velocity field is consistent with observations (Niederhofer et al., 2021; Zivick et al., 2021). Further, the SMC’s stellar velocity field in the scenario where the Clouds do not collide, remains a rotating disk (peak velocity $\sim 40$ km s-1). This is inconsistent with observations, indicating that strong LMC tides (like in a direct collision) are needed to sufficiently disrupt the SMC’s internal velocity field. The SMC’s rotation to dispersion ratio significantly reduces post-collision (<0.4<0.4, Figure 10), indicating the SMC is in the midst of a transformation from a rotationally supported disk to a dispersion dominated spheroid.
@@
    The SMC’s gas kinematics is dominated by radially outward motions as opposed to rotation: Pre-collision, the simulated SMC has a coherent gas rotation field, by design. Post-collision, the simulated SMC’s gas also exhibits radially outward motions and the rotation field is destroyed (Figure 11). The resulting diverging velocity field manifests as a velocity gradient (408040-80 km s-1) when viewed at an inclination. Hence, the SMC’s observed HI velocity gradient is likely due to radially outward motions of gas viewed at an inclination, and is not a signature of rotation. This finding is consistent with the kinematics of the SMC’s young stellar populations (<100<100 Myr, Murray et al., 2019; Nakano et al., 2025; Nakano and Tachihara, 2025). The SMC’s gas kinematics cannot be used to place the SMC on the BTFR (Figure 1 left panel).
@@
Our work highlights the crucial role of galactic collisions in transforming dIrr type galaxies to dE/dSph type over short timescales ($\sim 100$ Myr). Further, ram pressure exerted by the primary galaxy’s ISM during such close encounters can provide a significant velocity kick to the secondary’s gas, assisting the removal of gas and contributing to quenching of low mass galaxies. The SMC gives a front row view of group processes driving dramatic morphological and kinematic transformations, without requiring the presence of sustained environmental effects (like a cluster ICM). Such group processing may explain the how dEs can exist around massive galaxies like M31. Our work also suggests that caution must be exercised in interpreting the measured HI density profiles and velocity profiles of SMC like galaxies around MW like hosts in the local universe (Geha et al., 2024; Zhu et al., 2025), since those profiles can be significantly affected by tidal interactions in galaxy groups occurring prior to the group’s infall.
@@
Himansh Rathore would like to thank Annapurni Subramaniam, Smitha Subramanian, Claire Murray, Paul Zivick, Andrew Fox, Ivanna Escala, Nico Garavito Camargo, Kathryne J. Daniel, S.R. Dhanush, Mathieu Renzo, Yumi Choi, Dennis Zaritsky, Mariarosa Cioni, Lara Cullinane and Hayden Foote for interesting scientific discussions. We would like to thank the anonymous referee for providing insightful comments that improved the quality as well as clarity of the paper. Himansh Rathore and Gurtina Besla are supported by NASA FINESST 80NSSC24K1469, NASA ATP 80NSSC24K1225 and NSF CAREER AST 1941096. This work utilized the Puma and ElGato High Performance Computing clusters at the University of Arizona. The Theoretical Astrophysics Program (TAP) at the University of Arizona provided resources to support this work. We respectfully acknowledge the University of Arizona is on the land and territories of Indigenous peoples. Today, Arizona is home to 22 federally recognized tribes, with Tucson being home to the O’odham and the Yaqui. The university strives to build sustainable relationships with sovereign Native Nations and Indigenous communities through education offerings, partnerships, and community service.